Find average wage rate from the data given below:

Wage rate (Rs.)

Less than 20

20-40

40-60

60-80

More than 80

no. of workers 

4

6

5

3

2



1
Expert's answer
2021-07-01T13:18:01-0400

Class Mark = midpoint of an interval = (Lower limit+Upper limit) / 2

  1. (0 + 20) / 2 = 10
  2. (20 + 40) / 2 = 30
  3. (40 + 60) / 2 = 50
  4. (60 + 80) / 2 = 70
  5. (80 + 100) / 2 = 90


Average wage rate = (10 x 4) + (30 x 6) + (50 x 5) + (70 x 3) + (90 x 2) / 20 = 38